solve using cramer's rule:
4x-3y=-7
3x-9y=15
x=?, y=?
i put these equations in the augmented matrix form, but i was stuck on how to actually solve the equation. i thought that you were only able to solve cramers rule if you are given a square matrix.
can you please help me solve this equation?

Answer by Alan3354(30983)

solve using cramer's rule:
4x-3y=-7
3x-9y=15
-----------
Det = -36+9 = -27
------
For x:
-7 -3
15 -9
--> 63 + 45 = 108
x = -4
===============
For y:
4 -7
3 15
--> 60 + 21 = 81
x = -3